solve by elimination method
0.3x-0.2y=4
0.2x+0.3y=-37/17

Answer by checkley77(12844)

0.3x-0.2y=4 multiply by 1.5 & add.
0.2x+0.3y=-37/17
0.45x-0.3y=6
---------------------
0.65x=6-37/17
0.65x=(6*17-37)/17
0.65x=(102-37)/17
0.65x=65/17
x=(65/17)/0.65
x=(65/17)(1/0.65
x=5.88 ans.
--------------
0.3*5.88-0.2y=4
1.764-0.2y=4
-0.2y=4-1.764
-0.2y=2.236
y=2.236/-0.2
y=-11.18 ans.